OpenStudy (anonymous):

A motorboat travels 684 km in 9 hours going upstream and 378 km in three hours going downstream. Whatis the rate of the boat in still water and what is the rate of the current?

OpenStudy (kropot72):

@learning1 I remember trying to work with you on this question. You just had to solve a pair of simultaneous equations:(

OpenStudy (anonymous):

Let the rate of the boat=x and the rate of the stream =y. When the boat is going up the stream, it is working against the stream. So, its overall rate R=x-y. When it is traveling in the same direction as the stream, its overall rate R=x+y. Now, D=RT. Distance =Rate x Time. Distance traveled up the stream in T=9 hours is 684 km. Distance traveling down the stream in T=3 hours is 378. Now substituting all of our information together, we get: Upstream: 684=(x-y)9 and Downstream: 378=(x+y)3 Divide both sides of the first equation by 9: 76=x-y Divide both sides of the second equation by 3: 126=x+y Add the two equations together and the y's will cancel. This leaves us with: 202=2x. Divide both sides by 2: x=101 km/hr. This is the rate of your boat in still water.